OpenStudy (hybrik):

what is the area of the bounded region |x|+|4y|=16

OpenStudy (anonymous):

I think I would start by rearrange the eqn -> |x| = 16 - |4y|. and then sketch the four bounding curves for x,-x and y,-y.

OpenStudy (hybrik):

Can you so |4y|=-|x|+16 also?

OpenStudy (amity):

There are 4 straight lines 4y+x=16,x-4y=16,x-4y+16=0, x+4y+16=0

OpenStudy (amity):

There are 4 straight lines 4y+x=16,x-4y=16,x-4y+16=0, x+4y+16=0

OpenStudy (anonymous):

Yep, both are fine.

OpenStudy (hybrik):

amity, urs makes a rhombus, i cant do graphs

OpenStudy (hybrik):

@Redcan What you do next

OpenStudy (hybrik):

@freckles

OpenStudy (anonymous):

Lets start with x = 16+4y. We know its a line so lets fine two points. When y=0 the x = ?

OpenStudy (hybrik):

16, when x=0 y = 4

OpenStudy (hybrik):

(16*2)*(4*2)/2?

OpenStudy (hybrik):

thats what i think but continue on

OpenStudy (anonymous):

yep so the first line is through (0,16), and (4,0) since if x=0 --> y=4

OpenStudy (anonymous):



OpenStudy (amity):

Then 1st two lines intersect yand xaxis at (0,4) (16,0) (0,-4) .(16,0) common . Now 2nd two at (0,4) (0,-4) and (-16,0).now think that two lines pairs are perpendicular....hence it is a square of side length[ (16-0)^2+(0-4)^2] area is side square that is 272square unit

OpenStudy (amity):

Then 1st two lines intersect yand xaxis at (0,4) (16,0) (0,-4) .(16,0) common . Now 2nd two at (0,4) (0,-4) and (-16,0).now think that two lines pairs are perpendicular....hence it is a square of side length[ (16-0)^2+(0-4)^2] area is side square that is 272square unit

OpenStudy (anonymous):

Yep, we were only at the draw the picture stage

OpenStudy (hybrik):

it says the answer is 128?

OpenStudy (hybrik):

oh i get it now, the guy witht he 4 equations was right, graph it and u ave 16*2 and 4*2 as P and Q, PQ/2 = 128

OpenStudy (amity):

Yes it is not a square area=4(1/2•16•4=128 i have mistaken there

OpenStudy (amity):

Yes it is not a square area=4(1/2•16•4=128 i have mistaken there

OpenStudy (hybrik):

rhombus area = PQ/2, P = 32, Q=8, when u graph it u see side lengths, u were right